media%2Fb5c%2Fb5cf2091-1be2-45c8-9a81-29
media%2F3e0%2F3e023c67-5664-4593-90dc-82
media%2Fe0d%2Fe0daea96-4db7-4c50-9ec9-c5Please do part A , B and C . Zoom to see clearly. Thanks
Answer #1

a) center of mass = mx^2 = 1.8*(0.14)^2 = 0.03528 Kg-m^2

b) center of mass = ((m1m2)/(m1+m2))*x^2 = ((1.8*6.5)/*(1.8+6.5))*(0.775-0.14)^2 = 0.57 Kg- m^2

c) center of mass= remains the same as (b) as the distance between mass 1 and 2 is the same as in part two